# Characterizing rational homogeneous spaces via $\Csuperscript\C\C^{*}$-actions

### v1

@@ 6. Actions with an isolated extremal fixed point
By [25, Lemma 4.5] we have $\nu^{+}(Y_{1})=1$, hence the sink ${\mathcal{M}}_{y}\cap{\mathcal{U}}_{1}^{-}$ (which is non-empty by Corollary 4.12) is zero dimensional. Every point of ${\mathcal{M}}_{y}$ can be connected to the sink by a chain of closures of orbits, so ${\mathcal{M}}_{y}$ is rationally chain connected, hence rationally connected.
For the second part, note that the restriction to ${\mathcal{M}}_{y}$ of the line bundle $D=p^{*}(\mathop{\rm det}\nolimits(p_{*}q^{*}L))$ satisfies, by Lemma 6.6, the hypotheses of Theorem 3.9, which allows us to conclude that $\rho({\mathcal{M}}_{y})\leq 2$. ∎
@@
For every $x\in X$ variety ${\mathcal{M}}_{x}$ is a smooth rationally connected variety. Moreover, denoting by ${\mathop{\rm N_{1}}\nolimits}({\mathcal{M}}_{x},{\mathcal{U}})$ the image of $i_{*}:{\mathop{\rm N_{1}}\nolimits}({\mathcal{M}}_{x})\to{\mathop{\rm N_{1}}\nolimits}({\mathcal{U}})$ then
$\rho({\mathcal{M}}_{x})=\dim{\mathop{\rm N_{1}}\nolimits}({\mathcal{M}}_{x},{\mathcal{U}})=\rho({\mathcal{U}})-\rho(X)\leq 2.$
@@
By Lemma 6.8, the fiber ${\mathcal{M}}_{y}$ of $q$ over a point of $Y_{1}$ is rationally connected. Since rational connectedness is an open and closed property on smooth proper families, all the fibers of $q$ are rationally connected. Since $X$ is rationally connected, by [11, Corollary 1.3] also ${\mathcal{U}}$ is rationally connected. Now we can use the arguments in [15, Proposition 3.1], to show that, for every $x\in X$ we have
$\rho({\mathcal{M}}_{x})=\dim{\mathop{\rm N_{1}}\nolimits}({\mathcal{M}}_{x},{\mathcal{U}})=\rho({\mathcal{U}})-\rho(X).$
Then we conclude by noting that $\rho({\mathcal{M}}_{y})\leq 2$ for $y\in Y_{1}$, by Lemma 6.8. ∎
@@
Let us start by noting that Theorem 1.1 holds, by Proposition 3.7, in the case in which the bandwidth of the action on $(X,L)$ is equal to two. Then we may assume, without loss of generality, that the bandwidth of the action is at least $3$. In this case the proof will follow from the description of the varieties ${\mathcal{M}}_{x}$, for every $x\in X$. We will show that every ${\mathcal{M}}_{x}$ is isomorphic to ${\mathcal{M}}_{y}$, for $y\in Y_{1}$, and that ${\mathcal{M}}_{y}$ is either one of the rational homogeneous cases of Theorem 3.6 if $\rho({\mathcal{U}}/X)=1$, or a product of two projective spaces if $\rho({\mathcal{U}}/X)=2$.
In both cases we know that the variety ${\mathcal{M}}_{x}$ coincides at every point with the family of tangent directions to lines at a point of a irreducible Hermitian symmetric space (see Table 4). Therefore, the statement will follow from Theorem 2.11.
@@
By Corollary 6.9, for every $x\in X$ the variety ${\mathcal{M}}_{x}$ is rationally connected of Picard number one, hence a Fano manifold. In particular the divisor $D=p^{*}(\mathop{\rm det}\nolimits(p_{*}q^{*}L))$, which has positive degree on rational curves in ${\mathcal{M}}_{x}$ by Lemma 6.6, must be ample on ${\mathcal{M}}_{x}$.
Now we consider a point $y\in Y_{1}$, and claim that the variety ${\mathcal{M}}_{y}$ is a rational homogeneous manifold. As observed in Lemma 6.8 the ${\mathbb{C}}^{*}$-action on $({\mathcal{M}}_{y},D_{|{\mathcal{M}}_{y}})$ has isolated sink and, by Lemma 6.6, it has bandwidth two.
If the criticality of the action is one (i.e. if the action on ${\mathcal{M}}_{y}$ has no inner fixed point components) then ${\mathcal{M}}_{y}$ is a projective space by Proposition 3.3. More precisely, the polarized pair $({\mathcal{M}}_{y},D_{|{\mathcal{M}}_{y}})$ is isomorphic to $({\mathbb{P}}^{m},{\mathcal{O}}_{{\mathbb{P}}^{m}}(2))$.
If the criticality is two, then ${\mathcal{M}}_{y}$ is one of the varieties appearing in Theorem 3.6, so we only need to exclude the non-homogeneous cases.
The morphism $q:{\mathcal{U}}\to X$, which is smooth of relative dimension $m$, is the contraction of an extremal ray, generated by the class of a curve of degree one with respect to $D$; denote by $k$ the anticanonical degree of this class. By Lemma 6.6, if $y$ is an inner fixed point of the action on $X$, the anticanonical degree of a general orbit of the induced action on ${\mathcal{M}}_{y}$ is $2k$. For every inner fixed point component we then have, by Corollary 2.4
@@
Let us note that, if equality holds, then $r=3$ and $u_{r}^{-}=m$; in particular the action on ${\mathcal{U}}_{r}$ is trivial, hence, by Proposition 5.2, $\dim Y_{r}=0$. This case has been treated in [26, Theorem 8.10]. Moreover, if $u_{r}^{-}=m-1$, then, for $x\in Y_{r}$, the action on ${\mathcal{M}}_{x}$, which has criticality one, has divisorial sink; since $\rho({\mathcal{M}}_{x})=1$ the source of the action on ${\mathcal{M}}_{x}$ must be isolated, and ${\mathcal{M}}_{x}$ is a projective space by Proposition 3.3, a contradiction. We can then assume $12(2k-m)<m-$ 1, that is, $24k\leq 3m-$ 2. Listing the varieties in Theorem 3.6 (5) and (6), with the corresponding values of $m$ and $k$, we see that this inequality is never satisfied.
@@
We then conclude that ${\mathcal{M}}_{y}$ is a rational homogeneous manifold. Since rational homogeneous manifolds have no nontrivial infinitesimal deformations, it follows that ${\mathcal{M}}_{x}\simeq{\mathcal{M}}_{y}$ for a general $x\in X$. Since every possible case for ${\mathcal{M}}_{y}$ that we have is rigid under smooth Fano deformations (cf. [13, Main Theorem]), it follows that every ${\mathcal{M}}_{x}$ is isomorphic to ${\mathcal{M}}_{y}$, and we may conclude the proof by [27, Theorem 1.1].
@@
We consider again a point $y\in Y_{1}$. As observed in Lemma 6.8 the ${\mathbb{C}}^{*}$-action on $({\mathcal{M}}_{y},D_{|{\mathcal{M}}_{y}})$ has isolated sink and, by Lemma 6.6, it has bandwidth two. By Theorem 3.9 then either
@@
    there exists $0<a<m$ such that ${\mathcal{M}}_{y}\simeq{\mathbb{P}}^{a}\times{\mathbb{P}}^{m-a}$, or
@@
    there exists $0<a<m$ such that ${\mathcal{M}}_{y}$ is isomorphic to the blowup of ${\mathbb{P}}^{m}$ along a linear space of dimension $a-1$.
@@
We will now construct two unsplit families of rational curves on ${\mathcal{U}}$, starting from rational curves in ${\mathcal{M}}_{y}$. We denote by $\ell_{1}$ and $\ell_{2}$ minimal rational curves in the extremal rays of ${\mathcal{M}}_{y}$. In case (P) they are lines in the factors of ${\mathcal{M}}_{y}$, in case (B) we choose $\ell_{1}$ in the ray corresponding to the fiber-type contraction ${\mathcal{M}}_{y}\to{\mathbb{P}}^{m-a}$ and $\ell_{2}$ in the ray corresponding to the blowup. Note that, by Corollary 6.9, the classes of the curves $\ell_{1}$ and $\ell_{2}$ are numerically independent in ${\mathop{\rm N_{1}}\nolimits}({\mathcal{U}})$.
@@
Denote by $V^{1}$ and $V^{2}$ the components of $\mathop{\rm RatCurves}\nolimits^{n}({\mathcal{U}})$ containing the elements corresponding to $\ell_{1}$ and $\ell_{2}$. By Lemma 6.7, $V^{1}$ and $V^{2}$ are unsplit families. Moreover, since $\ell_{1}$ is a free rational curve in ${\mathcal{M}}_{y}$ and the normal bundle of ${\mathcal{M}}_{y}$ in ${\mathcal{U}}$ is trivial, then $\ell_{1}$ is a free rational curve in ${\mathcal{U}}$; in particular the family $V^{1}$ is a covering family.
@@
In case (P) a similar argument shows that also $V^{2}$ is an unsplit covering family. Using that, by adjunction, $-K_{{\mathcal{U}}}|_{{\mathcal{M}}_{y}}=-K_{{\mathcal{M}}_{y}}$, we can compute the anticanonical degree of curves parametrized by $V^{1}$ and $V^{2}$; those degrees are, respectively, $a+1$ and $m-a+1$. Let ${\mathcal{M}}_{x}$ be any fiber of $q$. Then the curves parametrized by $V^{i}$ meeting ${\mathcal{M}}_{x}$ are contained in ${\mathcal{M}}_{x}$, so ${\mathcal{M}}_{x}$ admits two unsplit covering families of rational curves such that the sum of their anticanonical degrees is $\dim{\mathcal{M}}_{x}+2$. We conclude that ${\mathcal{M}}_{x}\simeq{\mathbb{P}}^{a}\times{\mathbb{P}}^{m-a}$ by [22, Main Theorem], for every $x\in X$, and this implies that $X$ is a rational homogeneous variety by [27, Theorem 1.1].
@@
We will finish the proof by showing that case (B) cannot occur. We claim first that, in this case $q(\mathop{\rm Locus}\nolimits(V^{2}))=X$. To prove the claim, since $V^{2}$ is unsplit, it is enough to check that the evaluation of global sections
$H^{0}({T_{\mathcal{U}}}_{|\ell_{2}})\otimes{\mathcal{O}}_{\ell_{2}}\to{q^{*}T_{X}}_{|\ell_{2}}$
@@
is surjective. Since ${T_{{\mathcal{M}}_{y}}}_{|\ell_{2}}\simeq{\mathcal{O}}_{\ell_{2}}(2)\oplus{\mathcal{O}}_{\ell_{2}}(1)^{m-a-1}\oplus{\mathcal{O}}_{\ell_{2}}^{a-1}\oplus{\mathcal{O}}_{\ell_{2}}(-1)$, and $q^{*}{T_{X}}_{|\ell_{2}}$ is trivial, then we have a commutative diagram with exact rows of the form:
@@
It follows that, for every $x\in X$, the variety ${\mathcal{M}}_{x}$ admits two unsplit families of rational curves, $V_{x}^{1}$ and $V_{x}^{2}$, of degree one with respect to $D$ and such that $V_{x}^{1}$ is a covering family. In particular this is the case for $x=y_{0}$. Then, by Lemma 6.4, we see that $Y_{1}$ admits two unsplit families of rational curves of $L$-degree one, $W^{1}$ and $W^{2}$, such that $W^{1}$ is a covering family. By Corollary 6.5, curves parametrized by these two families belong to ${\mathcal{M}}$. In particular, for a point $y\in Y_{1}$ belonging to $\mathop{\rm Locus}\nolimits(W^{2})$, the subvariety of ${\mathcal{M}}$ consisting of curves contained in $Y_{1}$ and passing by $y$ has two disjoint components. This contradicts the fact that the subvariety of ${\mathcal{M}}$ consisting of curves contained in $Y_{1}$ and passing by $y$ is the union of the inner fixed point components of the action of ${\mathcal{M}}_{y}$, which is a linear space (see the Proof of Theorem 3.9).

### v2

@@ 6. Actions with an isolated extremal fixed point
By [23, Lemma 4.5] we have $\nu^{+}(Y_{1})=1$, hence the sink ${\mathcal{U}}_{y}\cap{\mathcal{U}}_{1}^{-}$ (which is non-empty by Corollary 4.12) is zero dimensional. Every point of ${\mathcal{U}}_{y}$ can be connected to the sink by a chain of closures of orbits, so ${\mathcal{U}}_{y}$ is rationally chain connected, hence rationally connected.
For the second part, note that the restriction to ${\mathcal{U}}_{y}$ of the line bundle $D=p^{*}(\mathop{\rm det}\nolimits(p_{*}q^{*}L))$ satisfies, by Lemma 6.6, the hypotheses of Theorem 3.9, which allows us to conclude that $\rho({\mathcal{U}}_{y})\leq 2$. ∎
@@
For every $x\in X$ the variety ${\mathcal{U}}_{x}$ is a smooth rationally connected variety. Moreover, denoting by ${\mathop{\rm N_{1}}\nolimits}({\mathcal{U}}_{x},{\mathcal{U}})$ the image of $i_{*}:{\mathop{\rm N_{1}}\nolimits}({\mathcal{U}}_{x})\to{\mathop{\rm N_{1}}\nolimits}({\mathcal{U}})$, we have:
$\rho({\mathcal{U}}_{x})=\dim{\mathop{\rm N_{1}}\nolimits}({\mathcal{U}}_{x},{\mathcal{U}}).$
In particular, if $\rho({\mathcal{U}}/X)=1$, then $\rho({\mathcal{U}}_{x})=1$ for every $x\in X$.
@@
By Lemma 6.8, the fiber ${\mathcal{U}}_{y}$ of $q$ over a point of $Y_{1}$ is rationally connected. Since rational connectedness is an open and closed property on smooth proper families, all the fibers of $q$ are rationally connected. Since $X$ is rationally connected, by [11, Corollary 1.3] also ${\mathcal{U}}$ is rationally connected. Now we follow the arguments of [15, Proposition 3.1]: for every $x\in X$ the map
$\mathop{\rm Pic}\nolimits({\mathcal{U}})\otimes{\mathbb{Q}}\simeq H^{2}({\mathcal{U}},{\mathbb{Q}})\to H^{2}({\mathcal{U}}_{x},{\mathbb{Q}})\simeq\mathop{\rm Pic}\nolimits({\mathcal{U}}_{x})\otimes{\mathbb{Q}}$
is surjective by Deligne’s Invariant Cycle Theorem. In particular, for every $x\in X$ we have $\rho({\mathcal{U}}_{x})=\dim N_{1}({\mathcal{U}}_{x},{\mathcal{U}})$. If $\rho({\mathcal{U}}/X)=1$, then $q$ is a Mori contraction of fiber type, hence $\dim N_{1}({\mathcal{U}}_{x},{\mathcal{U}})\leq\rho({\mathcal{U}}/X)$, and so $\rho({\mathcal{U}}_{x})=1$. ∎
@@
Let us start by noting that Theorem 1.1 holds, by Proposition 3.7, in the case in which the bandwidth of the action on $(X,L)$ is equal to two. Then we may assume, without loss of generality, that the bandwidth of the action is at least $3$. In this case the proof will follow from the description of the varieties ${\mathcal{U}}_{x}$, for every $x\in X$. We will show that every ${\mathcal{U}}_{x}$ is isomorphic to ${\mathcal{U}}_{y}$, for $y\in Y_{1}$, and that ${\mathcal{U}}_{y}$ is either one of the rational homogeneous cases of Theorem 3.6 if $\rho({\mathcal{U}}/X)=1$, or a product of two projective spaces if $\rho({\mathcal{U}}/X)\geq 2$.
In both cases we know that the variety ${\mathcal{U}}_{x}$ coincides at every point with the family of tangent directions to lines at a point of an irreducible Hermitian symmetric space (see Table 4). Hence, the statement will follow from Theorem 2.12.
@@
By Corollary 6.9, for every $x\in X$ the variety ${\mathcal{U}}_{x}$ is rationally connected of Picard number one, hence a Fano manifold. In particular the divisor $D=p^{*}(\mathop{\rm det}\nolimits(p_{*}q^{*}L))$, which has positive degree on rational curves in ${\mathcal{U}}_{x}$ by Lemma 6.6, must be ample on ${\mathcal{U}}_{x}$.
Now we consider a point $y\in Y_{1}$, and claim that the variety ${\mathcal{U}}_{y}$ is a rational homogeneous manifold. As observed in Lemma 6.8 the ${\mathbb{C}}^{*}$-action on $({\mathcal{U}}_{y},D_{|{\mathcal{U}}_{y}})$ has isolated sink and, by Lemma 6.6, it has bandwidth two.
If the criticality of the action is one (i.e. if the action on ${\mathcal{U}}_{y}$ has no inner fixed point components) then ${\mathcal{U}}_{y}$ is a projective space by Proposition 3.3. More precisely, the polarized pair $({\mathcal{U}}_{y},D_{|{\mathcal{U}}_{y}})$ is isomorphic to $({\mathbb{P}}^{m},{\mathcal{O}}_{{\mathbb{P}}^{m}}(2))$.
If the criticality is two, then ${\mathcal{U}}_{y}$ is one of the varieties appearing in Theorem 3.6, so we only need to exclude the non-homogeneous cases.
The morphism $q:{\mathcal{U}}\to X$, which is smooth of relative dimension $m$, is the contraction of an extremal ray, generated by the class of a curve of degree one with respect to $D$; denote by $k$ the anticanonical degree of this class. By Lemma 6.6, if $y$ is an inner fixed point of the action on $X$, the anticanonical degree of a general orbit of the induced action on ${\mathcal{U}}_{y}$ is $2k$. Then, by Corollary 2.4, for every inner fixed point component we have:
@@
Let us note that, if equality holds, then $r=3$ and $u_{r}^{-}=m$; in particular the action on ${\mathcal{U}}_{r}$ is trivial, hence, by Proposition 5.2, $\dim Y_{r}=0$. This case has been treated in [26, Theorem 8.10]. Moreover, if $u_{r}^{-}=m-1$, then, for $x\in Y_{r}$, the action on ${\mathcal{U}}_{x}$, which has criticality one, has divisorial sink; since $\rho({\mathcal{U}}_{x})=1$ the source of the action on ${\mathcal{U}}_{x}$ must be isolated, and ${\mathcal{U}}_{x}$ is a projective space by Proposition 3.3, a contradiction. We can then assume $12(2k-m)<m-$ 1, that is, $24k\leq 3m-$ 2. Listing the varieties in Theorem 3.6 (5) and (6), with the corresponding values of $m$ and $k$, we see that this inequality is never satisfied.
@@
We then conclude that ${\mathcal{U}}_{y}$ is a rational homogeneous manifold. Since rational homogeneous manifolds have no non-trivial infinitesimal deformations, it follows that ${\mathcal{U}}_{x}\simeq{\mathcal{U}}_{y}$ for a general $x\in X$. Since every possible case for ${\mathcal{U}}_{y}$ that we have is rigid under smooth Fano deformations (cf. [13, Main Theorem]), it follows that every ${\mathcal{U}}_{x}$ is isomorphic to ${\mathcal{U}}_{y}$, and we may conclude the proof by [27, Theorem 1.1].
@@
We consider again a point $y\in Y_{1}$. As observed in Lemma 6.8 the ${\mathbb{C}}^{*}$-action on $({\mathcal{U}}_{y},D_{|{\mathcal{U}}_{y}})$ has isolated sink and, by Lemma 6.6, it has bandwidth two. Then, by Theorem 3.9, either
@@
    there exists $0<a<m$ such that ${\mathcal{U}}_{y}\simeq{\mathbb{P}}^{a}\times{\mathbb{P}}^{m-a}$, or
@@
    there exists $0<a<m$ such that ${\mathcal{U}}_{y}$ is isomorphic to the blowup of ${\mathbb{P}}^{m}$ along a linear space of dimension $a-1$.
@@
We will now construct two unsplit families of rational curves on ${\mathcal{U}}$, starting from rational curves in ${\mathcal{U}}_{y}$. We denote by $\ell_{1}$ and $\ell_{2}$ minimal rational curves in the extremal rays of ${\mathcal{U}}_{y}$. In case (P) they are lines in the factors of ${\mathcal{U}}_{y}$; in case (B) we choose $\ell_{1}$ in the ray corresponding to the fiber-type contraction ${\mathcal{U}}_{y}\to{\mathbb{P}}^{m-a}$ and $\ell_{2}$ in the ray contracted by the blowup map. Note that, by Corollary 6.9, the classes of the curves $\ell_{1}$ and $\ell_{2}$ are numerically independent in ${\mathop{\rm N_{1}}\nolimits}({\mathcal{U}})$.
@@
Denote by $V^{1}$ and $V^{2}$ the components of $\mathop{\rm RatCurves}\nolimits^{n}({\mathcal{U}})$ containing the elements corresponding to $\ell_{1}$ and $\ell_{2}$. By Lemma 6.7, $V^{1}$ and $V^{2}$ are unsplit families. Moreover, since $\ell_{1}$ is a free rational curve in ${\mathcal{U}}_{y}$ and the normal bundle of ${\mathcal{U}}_{y}$ in ${\mathcal{U}}$ is trivial, then $\ell_{1}$ is a free rational curve in ${\mathcal{U}}$; in particular the family $V^{1}$ is a covering family.
@@
In case (P) a similar argument shows that also $V^{2}$ is an unsplit covering family. Using that, by adjunction, $-K_{{\mathcal{U}}}|_{{\mathcal{U}}_{y}}=-K_{{\mathcal{U}}_{y}}$, we can compute the anticanonical degree of curves parametrized by $V^{1}$ and $V^{2}$; those degrees are, respectively, $a+1$ and $m-a+1$. Let ${\mathcal{U}}_{x}$ be any fiber of $q$. Then the curves parametrized by $V^{i}$ meeting ${\mathcal{U}}_{x}$ are contained in ${\mathcal{U}}_{x}$, so ${\mathcal{U}}_{x}$ admits two unsplit covering families of rational curves such that the sum of their anticanonical degrees is $\dim{\mathcal{U}}_{x}+2$. We conclude that ${\mathcal{U}}_{x}\simeq{\mathbb{P}}^{a}\times{\mathbb{P}}^{m-a}$ by [22, Main Theorem], for every $x\in X$, and this implies that $X$ is a rational homogeneous variety by [27, Theorem 1.1].
@@
We will finish the proof by showing that case (B) cannot occur. We claim first that, in this case, $q(\mathop{\rm Locus}\nolimits(V^{2}))=X$. To prove the claim, since $V^{2}$ is unsplit, it is enough to check that $q|_{\mathop{\rm Locus}\nolimits(V^{2})}$ has surjective differential at a point of $\ell_{2}$; we claim first that this is equivalent to saying that the evaluation of global sections
$H^{0}({T_{\mathcal{U}}}_{|\ell_{2}})\otimes{\mathcal{O}}_{\ell_{2}}\to{q^{*}T_{X}}_{|\ell_{2}}$ (14)
is surjective. In fact, as a component of $\mathop{\rm RatCurves}\nolimits^{n}({\mathcal{U}})$, $V^{2}$ is the normalization of the image of a subscheme of $\mathop{\rm Hom}\nolimits({\mathbb{P}}^{1},{\mathcal{U}})$, that we denote by ${\mathcal{H}}_{2}$. Our claim follows from proving that the composition ${\mathbb{P}}^{1}\times{\mathcal{H}}_{2}\to{\mathcal{U}}\to X$ has surjective differential at a point $(P,\ell_{2})$, for some $P\in\ell_{2}$ or, equivalently, that the composition
${\mathbb{P}}^{1}\times{\mathcal{H}}_{2}\subset{\mathbb{P}}^{1}\times\mathop{\rm Hom}\nolimits({\mathbb{P}}^{1},{\mathcal{U}})\to\mathop{\rm Hom}\nolimits({\mathbb{P}}^{1},X)\to X$
has surjective differential at $(P,\ell_{2})$; this follows from [17, II Proposition 3.4].
Let us now prove the surjectivity of the map in (14). Since
${T_{{\mathcal{U}}_{y}}}_{|\ell_{2}}\simeq{\mathcal{O}}_{\ell_{2}}(2)\oplus{\mathcal{O}}_{\ell_{2}}(1)^{m-a-1}\oplus{\mathcal{O}}_{\ell_{2}}^{a-1}\oplus{\mathcal{O}}_{\ell_{2}}(-1),$
@@
and $q^{*}{T_{X}}_{|\ell_{2}}$ is trivial, then we have a commutative diagram with exact rows of the form:
@@
It follows that, for every $x\in X$, the variety ${\mathcal{U}}_{x}$ admits two unsplit families of rational curves, $V_{x}^{1}$ and $V_{x}^{2}$, of degree one with respect to $D$ and such that $V_{x}^{1}$ is a covering family. In particular, this is the case for $x=y_{0}$. Then, by Lemma 6.4, we see that $Y_{1}$ admits two unsplit families of rational curves of $L$-degree one, $W^{1}$ and $W^{2}$, such that $W^{1}$ is a covering family. By Corollary 6.5, curves parametrized by these two families belong to ${\mathcal{M}}$. In particular, for a point $y\in Y_{1}$ belonging to $\mathop{\rm Locus}\nolimits(W^{2})$, ${\mathcal{U}}_{y}^{0}={\mathcal{U}}_{Y_{1}}^{0}\cap{\mathcal{U}}_{y}$ has two disjoint components. This contradicts the fact that, by assumption, ${\mathcal{U}}_{y}$ is ${\mathbb{C}}^{*}$-equivariantly isomorphic to the blowup of a projective space endowed with a ${\mathbb{C}}^{*}$-action that has a unique inner fixed point component (see the Proof of Theorem 3.9).